# Evaluating Membership Inference Attacks in heterogeneous-data setups

### v1

@@ 3.2 Results
Figure 2: Two non-members sampling techniques: Humphries et al. [13] and ours.
Non-members Average Accuracy
0% from 3rd party (Our naive method) 47.75
25% from 3rd party 60.71
50% from 3rd party 57.78
75% from 3rd party 70.24
100% from 3rd party (Humphries et al. [13]) 91.23
@@
Table 2 presents the MIA accuracy for a varying proportion of third-party non-members: 0% corresponds to our sampling, and 100% corresponds to [13]. This table confirms that the more non-members are sampled from a third-party distribution, the higher the MIA accuracy is. With 100%, we obtain results similar to those reported in [13]. Thus, the non-member sampling was the cause of the contradictory results.
@@ 4.0.1 Acknowledgements
This work is based on the MSc thesis of Bram van Dartel [5].

### v2

@@ 3.2 Results
Figure 2: Approaches to sample heterogeneous data in MIA
Sampling Average Accuracy
2-distribution setting 47.75
25% of non-members from 3rd distribution 60.71
50% of non-members from 3rd distribution 57.78
75% of non-members from 3rd distribution 70.24
3-distribution seeting (like in [13]) 91.23
@@
Table 2 presents the MIA accuracy for a varying proportion of non-members drawn from a third distribution: 0% corresponds to our 2-distribution setting, and 100% corresponds to the 3-distribution setting of [13]. This table confirms that the more non-members are sampled from a third-party distribution, the higher the MIA accuracy is. With 3 distributions, we obtain results similar to those reported in [13]. Thus, the non-member sampling was the cause of the contradictory results.
@@ 4.0.1 Acknowledgements
This work is based on the MSc thesis of Bram van Dartel [5], and was supported by the Netherlands Organization for Scientific Research (De Nederlandse Organisatie voor Wetenschappelijk Onderzoek) under NWO:SHARE project [CS.011].